OpenStudy (anonymous):

Solve |8f+3|=5 for f. Show all work.

OpenStudy (dtan5457):

when you have a absolute value equation like that, it means 8f+3=5 8f+3=-5 now just solve for f, and you will end up with 2 values

OpenStudy (anonymous):

So, 8f = -8 (so f = -1)?

OpenStudy (dtan5457):

for the 2nd one, yes, f=-1 now just solve the first one

OpenStudy (anonymous):

What is the first one?

OpenStudy (dtan5457):

well if you can solve the 2nd one, use the same concepts for the first one

OpenStudy (anonymous):

I'm not sure how to solve for another answer?

OpenStudy (dtan5457):

8f+3=5 -3 -3 8f=2 f=?

OpenStudy (anonymous):

Oh. F = 4

OpenStudy (anonymous):

now i got it. thanks!

OpenStudy (dtan5457):

hold on, it's not 4, lol. 8f=2 im gonna divide everything by 8 so i can f alone f=2/8

OpenStudy (anonymous):

oopsss 1/4

OpenStudy (dtan5457):

or 0.25, same thing :)

OpenStudy (anonymous):

wow... going back to algebra 1 after algebra 2 and geometry is just hard.

OpenStudy (anonymous):

thanks